# Large Language Models and Code Security: A Systematic Literature Review

### v2

@@ IV-C File Management
File management vulnerabilities are mostly caused by improper handling of files and directories. These vulnerabilities can lead to unauthorized access, data corruption, or system compromise. The most common file management vulnerabilities include improper path restrictions, unrestricted file uploads, and insecure file permissions. Several studies have identified file management vulnerabilities in LLM-generated code [23, 24, 26, 28, 31, 33, 34, 38, 38, 39, 40]. However, He et al. [11], and Mohsin et al.[10] dedicated more focus to these vulnerabilities, finding that LLMs often produce code that does not properly restrict file paths based on user input, potentially allowing unauthorized file access. This vulnerability is marked as CWE-22: Improper Limitation of a Pathname to a Restricted Directory (’Path Traversal’) [23, 28, 31, 33, 38, 39, 10, 40].
IV-D Deserialization
@@
Deserialization vulnerabilities happen when untrusted data is used to reconstruct objects within an application. These vulnerabilities can lead to arbitrary code execution, data manipulation, or denial-of-service attacks. These vulnerabilities are especially concerning in systems that use serialization for communication or storage, as they can be exploited if proper validation and sanitization are not implemented. The primary CWE associated with this issue is: CWE-502: Deserialization of Untrusted Data [23, 33, 37, 39, 40].
@@
Several studies have recognized deserialization vulnerabilities [23, 33, 35, 37, 39, 40], however, none of them specifically focused on this topic, leaving some open questions for further research. For example, how frequently do LLMs generate insecure code when prompted with tasks targeting deserialization vulnerabilities?
IV-E Sensitive Data Exposure
@@
Several studies have recognized the exposure of sensitive data in the code generated by LLMs [23, 24, 29, 30, 38, 39, 10, 40]. However, Pearce et al. [28] dedicated more attention to this category of vulnerabilities. They studied how frequently and why GitHub Copilot generates insecure code by testing it with 89 different scenarios related to high-risk cybersecurity vulnerabilities. Besides from finding that approximately 40% of these programs were vulnerable, they showed that Copilot consistently fails when it comes to CWE-200: Exposure of Sensitive Information to an Unauthorized Actor. The same CWE was also mentioned in [23, 39, 10, 40]. A concrete example of how Copilot-generated code exhibits a CWE-200 vulnerability can be found with Python web apps [28]. In some cases, although the app is supposed to display only the last 4 digits of a Social Security Number (SSN), it happened that it revealed more digits than expected.
IV-F Authentication and Authorization
Authentication and authorization vulnerabilities represent a significant category of security issues, highlighted in 12 out of 20 studies.
Authentication vulnerabilities occur when a system fails to properly verify the identity of users or devices. This kind of vulnerability is mostly caused by weak password requirements, insufficiently protected credentials, or the use of hard-coded credentials.
@@ IV-F Authentication and Authorization
Hamer et al. [36] compared the security properties of the code generated by ChatGPT, with answers provided in Stack Overflow. They concluded that ChatGPT-generated code had 20% fewer vulnerabilities than Stack Overflow snippets, although both produced security flaws. One of the most frequent CWES in both ChatGPT-generated code and Stack Overflow snippets was CWE-798: Use of Hard-coded Credentials. CWE-798 is also discussed in [23, 28, 30, 40].
@@
Authorization vulnerabilities arise when a system fails to properly enforce access control policies, which allows users or processes to perform actions beyond their intended permissions. This vulnerability is often caused by improper implementation of access control. The associated CWE is CWE-284: Improper Access Control, as discussed by Asare et al. [25], who conducted a comparative empirical analysis of Copilot-generated code, showing that GitHub Copilot is less likely to produce vulnerable code, compared to human developers.
IV-G Cryptography
@@
Cryptographic vulnerabilities arise when encryption mechanisms are poorly implemented or completely absent. Common causes of weak encryption include the use of broken or insecure cryptographic algorithms, improper validation of certificates, the use of weak or absent hashing functions, and the reuse or hard-coding of cryptographic keys. Several studies [24, 30, 33, 36, 37, 38, 39, 10, 40] identified cryptographic weaknesses in LLM-generated code, without exploring them in detail.
@@
Perry et al. [26] conducted a user study to examine how users interact with AI code assistants to solve security-related tasks. The study found that participants with access to an AI assistant were more likely to produce incorrect and insecure solutions, compared to a control group. Specifically, in one cryptography-related task, participants with AI assistance were significantly more likely to use trivial ciphers or fail to authenticate the final output.
@@
A common cryptographic vulnerability identified in multiple studies is CWE-327: Use of a Broken or Risky Cryptographic Algorithm [33, 36, 37, 39].
IV-H Resource Management
@@
Hamer et al. [36] found that two CWEs directly related to resource management were particularly frequent in ChatGPT. CWE-404: Improper Resource Shutdown or Release, and CWE-772: Missing Release of Resource after Effective Lifetime. Both CWEs were tied as the second most frequent vulnerabilities, each appearing in 37, out of 216 code snippets. Other CWEs within the same vulnerability category were also highlighted in the studies, including CWE-605: Multiple Binds to the Same Port [37, 40], and CWE-664: Improper Control of a Resource Through its Lifetime [25].
In general, resource management vulnerabilities were identified by [25, 27, 32, 37, 38, 10, 40], however, none of these studies explored in depth the resource management vulnerabilities in LLM generated-code.
IV-I Coding Standards
@@
Tihanyi et al. [35] conducted a comparative analysis of LLMs to examine how likely they are to generate vulnerabilities when writing simple C programs. They identified division by zero as a frequent issue in the code produced by LLMs, highlighting the difficulties in managing arrays and accurately performing arithmetic operations.
@@
Additionally, CWE-758: Reliance on Undefined, Unspecified, or Implementation-Defined Behavior is another significant concern discussed in by Kim et al. [32], and Sandoval et al. [27]. Besides the mentioned studies, others have focused on this type of vulnerability, but without exploring it in detail, as can be seen in [25, 29, 38, 40].
IV-J Error Handling
@@
Error handling vulnerabilities arise when a system or application fails to properly manage exceptional conditions. Poor error handling can lead to a variety of issues, including security vulnerabilities, system crashes, or unreliable software behavior. One common issue in error handling is the failure to check or handle return values from functions or system calls properly. A few studies focused on error handling vulnerabilities [27, 29, 37], however, none of these studies specifically addressed this category of vulnerabilities, which presents another topic for future research. Among the various CWS, CWE-703: Improper Check or Handling of Exceptional Conditions has been specifically highlighted in the literature as a notable issue [37].
IV-K LLMs in Code Generation: Balancing Risks and Benefits
@@
As we have described throughout this section, LLMs often produce code with security vulnerabilities that developers might overlook [26, 29, 30, 34, 35, 38]. Interestingly, when ChatGPT fixes code initially written by humans, the resulting code often contains more security vulnerabilities compared to code generated from scratch by ChatGPT [37].
@@
Despite these challenges, LLMs remain valuable tools in coding tasks. Research indicates that code generated by ChatGPT generally introduces fewer CWE issues compared to code found on Stack Overflow [36]. Moreover, LLMs do not significantly increase the incidence rates of severe security bugs [27], and tools like GitHub Copilot have been shown to produce code with fewer security vulnerabilities than code written by human developers [25].
@@
Even though LLMs may generate vulnerable code, the literature suggests that using more effective prompts can guide them to produce safer code [23, 24]. Through improved re-prompting, the security issues initially present in LLM-generated code can often be addressed, making LLMs valuable tools for creating secure code, when used appropriately. This topic will be explored more in details in Section VII.
V LLMs for Vulnerability Detection
@@
Security code reviews performed at the early stages of software development can prevent the introduction of security flaws into code repositories. This helps to reduce future costs, associated with fixing and maintaining the software. Security code reviews performed manually require significant time and effort, especially in large-scale open-source projects with large numbers of contributions. Hence, the automated tools able to identify security vulnerabilities during code review are highly beneficial [43]. Among these tools, LLMs emerged as a promising solution for detecting security vulnerabilities.
@@
V-A Uses of LLMs for Vulnerability Detection
@@
The work by Mohajer et al. [44] demonstrated the ability of ChatGPT to perform static bug detection and false positive warning removal. The evaluation focused on two critical types of bugs often targeted by Static Analysis Tools (SATs): Null Dereference and Resource Leak. For detecting Null Dereference, bugsChatGPT achieved around 68% accuracy and 64% precision, and for detecting Resource Leak bugs, about 77% accuracy and 83% precision. Concerning false-positive warning removal, ChatGPT reached around 94% precision for Null Dereference bugs and 63% for Resource Leak bugs.
@@
Thapa et al. [45] empirically evaluated the performance of several transformer-based language models for software vulnerability detection. The authors presented a framework that allows for translating C/C++ source code into vectorized inputs, a format suitable for vulnerability analysis. The authors tested various models on software vulnerability datasets, and found that, concerning Buffer Errors and Resource Management Errors, GPT-2 Large and GPT-2 XL achieved the highest F1-score of all the models, with 95.51% and 95.40%, respectively.
@@
A study by Zhang et al. [46] focused on LLMs capabilities for automated vulnerability detection. They tested over 10 different LLM models using various techniques like zero-shot and one-shot learning (for more details on prompting techniques, we refer the reader to Section VIII), and fine-tuning. They applied these models to various datasets of source code, exhibiting C/C++ and smart contract vulnerabilities. The study found that CodeLlama-7B achieved the highest F1-score of 82% with discriminative fine-tuning, along with a precision of approximately 89% and a recall of around 78%. In comparison, other methods were less effective.
@@
According to Wen et al. [47], existing pre-trained models for vulnerability detection often miss important details and struggle with complex code. As a consequence, performing accurate vulnerability detection can be difficult. To address these challenges, Wen et al. [47] proposed a Structured Natural Language Comment tree-based Vulnerability Detection framework, based on pre-trained models, named SCALE. This framework is composed of three main modules. The first module, Comment Tree Construction, helps the model understand code semantics, by incorporating LLMs for comment generation and adding the comment node to Abstract Syntax Trees (ASTs). The second module, Structured Natural Language Comment Tree Construction, refines this process by explicitly involving code execution sequences and combining code syntax templates with the comment tree. The third module, SCT-Enhanced Representation, uses these improved comment trees to capture vulnerability patterns and better understand vulnerabilities in the code. Results show that SCALE outperforms existing methods, including supervised, retrained model-based, and other LLM-based approaches in detecting vulnerabilities.
@@
Omar et al. [48] proposed a Knowledge Distillation (KD) technique, designed to improve vulnerability detection. KD is a technique where a smaller, simpler model (the “student”) learns to replicate the behavior of a larger, more complex model (the “teacher”). The goal is to achieve similar or even better performance with the smaller model [49]. They assessed the performance of the KD technique when applied to various classifiers, including LLMs like GPT-2 and CodeBERT, as well as Recurrent Neural Network (RNN) models such as LSTM. Their evaluation revealed that GPT-2 performed outstandingly, achieving an F1 score of 92.4% on the SARD dataset [50]. While GPT-2 and CodeBERT showcased promising results, it was evident that LSTM-based models still have certain limitations in the context of this application.
@@
Wen et al. [51] highlighted that current methods and datasets focus on detecting intra-procedural vulnerabilities, those found within single functions, overlooking vulnerabilities occurring across multiple functions. To address this, they proposed VulEval, a repository-level evaluation system designed to evaluate both inter- and intra-procedural vulnerabilities. VulEval consists of three phases: the first detects intra-procedural vulnerabilities from code, the second retrieves the most relevant dependencies from call graphs, and the third detects inter-procedural vulnerabilities by incorporating these dependencies. Using VulEval over a large-scale dataset, which includes both intra- and inter-procedural vulnerabilities, they evaluated 19 detection methods: 4 program analysis-based methods, 2 supervised learning-based methods, 7 fine-tuning-based methods, and 6 prompt-based methods. Overall, incorporating vulnerability-related dependencies improved repository-level detection performance, over function-level detection.
V-B Issues in Using LLMs for Vulnerability Detection

### v3

@@ 4.3. File Management
File management vulnerabilities are mostly caused by improper handling of files and directories. These vulnerabilities can lead to unauthorized access, data corruption, or system compromise. The most common file management vulnerabilities include improper path restrictions, unrestricted file uploads, and insecure file permissions. Several studies have identified file management vulnerabilities in LLM-generated code (Liu et al., 2024c; Khoury et al., 2023; Perry et al., 2023; Pearce et al., 2022; Hajipour et al., 2024; Siddiq and Santos, 2022; Tóth et al., 2024; Fu et al., 2023, 2023; He et al., 2024; Gong et al., 2024). However, He et al. (He and Vechev, 2023), and Mohsin et al.(Mohsin et al., 2024) dedicated more focus to these vulnerabilities, finding that LLMs often produce code that does not properly restrict file paths based on user input, potentially allowing unauthorized file access. This vulnerability is marked as CWE-22: Improper Limitation of a Pathname to a Restricted Directory (’Path Traversal’) (Liu et al., 2024c; Pearce et al., 2022; Hajipour et al., 2024; Siddiq and Santos, 2022; Fu et al., 2023; He et al., 2024; Mohsin et al., 2024; Gong et al., 2024).
4.4. Deserialization
@@
Deserialization vulnerabilities happen when untrusted data is used to reconstruct objects within an application. These vulnerabilities can lead to arbitrary code execution, data manipulation, or denial-of-service attacks. These vulnerabilities are especially concerning in systems that use serialization for communication or storage, as they can be exploited if proper validation and sanitization are not implemented. The primary CWE associated with this issue is: CWE-502: Deserialization of Untrusted Data (Liu et al., 2024c; Siddiq and Santos, 2022; Rabbi et al., 2024; He et al., 2024; Gong et al., 2024).
@@
Several studies have recognized deserialization vulnerabilities (Liu et al., 2024c; Siddiq and Santos, 2022; Tihanyi et al., 2024; Rabbi et al., 2024; He et al., 2024; Gong et al., 2024), however, none of them specifically focused on this topic, leaving some open questions for further research. For example, how frequently do LLMs generate insecure code when prompted with tasks targeting deserialization vulnerabilities?
4.5. Sensitive Data Exposure
@@
Several studies have recognized the exposure of sensitive data in the code generated by LLMs (Liu et al., 2024c; Khoury et al., 2023; Elgedawy et al., 2024; Siddiq et al., 2024; Fu et al., 2023; He et al., 2024; Mohsin et al., 2024; Gong et al., 2024). However, Pearce et al. (Pearce et al., 2022) dedicated more attention to this category of vulnerabilities. They studied how frequently and why GitHub Copilot generates insecure code by testing it with 89 different scenarios related to high-risk cybersecurity vulnerabilities. Besides from finding that approximately 40% of these programs were vulnerable, they showed that Copilot consistently fails when it comes to CWE-200: Exposure of Sensitive Information to an Unauthorized Actor. The same CWE was also mentioned in (Liu et al., 2024c; He et al., 2024; Mohsin et al., 2024; Gong et al., 2024). A concrete example of how Copilot-generated code exhibits a CWE-200 vulnerability can be found with Python web apps (Pearce et al., 2022). In some cases, although the app is supposed to display only the last 4 digits of a Social Security Number (SSN), it happened that it revealed more digits than expected.
4.6. Authentication and Authorization
@@
Authentication and authorization vulnerabilities represent a significant category of security issues, highlighted in 12 out of 20 studies. Authentication vulnerabilities occur when a system fails to properly verify the identity of users or devices. This kind of vulnerability is mostly caused by weak password requirements, insufficiently protected credentials, or the use of hard-coded credentials. Hamer et al. (Hamer et al., 2024) compared the security properties of the code generated by ChatGPT, with answers provided in Stack Overflow. They concluded that ChatGPT-generated code had 20% fewer vulnerabilities than Stack Overflow snippets, although both produced security flaws. One of the most frequent CWES in both ChatGPT-generated code and Stack Overflow snippets was CWE-798: Use of Hard-coded Credentials. CWE-798 is also discussed in (Liu et al., 2024c; Pearce et al., 2022; Siddiq et al., 2024; Gong et al., 2024).
@@
Authorization vulnerabilities arise when a system fails to properly enforce access control policies, which allows users or processes to perform actions beyond their intended permissions. This vulnerability is often caused by improper implementation of access control. The associated CWE is CWE-284: Improper Access Control, as discussed by Asare et al. (Asare et al., 2023), who conducted a comparative empirical analysis of Copilot-generated code, showing that GitHub Copilot is less likely to produce vulnerable code, compared to human developers.
4.7. Cryptography
@@
Cryptographic vulnerabilities arise when encryption mechanisms are poorly implemented or completely absent. Common causes of weak encryption include the use of broken or insecure cryptographic algorithms, improper validation of certificates, the use of weak or absent hashing functions, and the reuse or hard-coding of cryptographic keys. Several studies (Khoury et al., 2023; Siddiq et al., 2024; Siddiq and Santos, 2022; Hamer et al., 2024; Rabbi et al., 2024; Fu et al., 2023; He et al., 2024; Mohsin et al., 2024; Gong et al., 2024) identified cryptographic weaknesses in LLM-generated code, without exploring them in detail.
@@
Perry et al. (Perry et al., 2023) studied how users interact with AI code assistants to solve security-related tasks. The study found that participants with access to an AI assistant were more likely to produce incorrect and insecure solutions, compared to a control group. Specifically, in one cryptography-related task, participants with AI assistance were significantly more likely to use trivial ciphers or fail to authenticate the final output.
@@
A common cryptographic vulnerability identified in multiple studies is CWE-327: Use of a Broken or Risky Cryptographic Algorithm (Siddiq and Santos, 2022; Hamer et al., 2024; Rabbi et al., 2024; He et al., 2024).
4.8. Resource Management
@@
Hamer et al. (Hamer et al., 2024) identified two prevalent resource management vulnerabilities in ChatGPT-generated code: CWE-404: Improper Resource Shutdown or Release, and CWE-772: Missing Release of Resource after Effective Lifetime. Both CWEs were tied as the second most frequent vulnerabilities, each appearing in 37, out of 216 code snippets. Other studies also highlighted related vulnerabilities, including CWE-605: Multiple Binds to the Same Port (Rabbi et al., 2024; Gong et al., 2024), and CWE-664: Improper Control of a Resource Through its Lifetime (Asare et al., 2023).
In general, resource management vulnerabilities were identified by (Asare et al., 2023; Sandoval et al., 2023; Kim et al., 2024; Rabbi et al., 2024; Fu et al., 2023; Mohsin et al., 2024; Gong et al., 2024), however, none of these studies explored in depth the resource management vulnerabilities in LLM generated-code.
4.9. Coding Standards
@@
Tihanyi et al. (Tihanyi et al., 2024) conducted a comparative analysis of LLMs to examine how likely they are to generate vulnerabilities when writing simple C programs. They identified division by zero as a frequent issue in the code produced by LLMs, along with the challanges in managing arrays and arithmetic operations.
@@
Additionally, CWE-758: Reliance on Undefined, Unspecified, or Implementation-Defined Behavior is another significant concern discussed in by Kim et al. (Kim et al., 2024), and Sandoval et al. (Sandoval et al., 2023). Besides the mentioned studies, others have focused on this type of vulnerability, but without exploring it in detail, as can be seen in (Asare et al., 2023; Elgedawy et al., 2024; Fu et al., 2023; Gong et al., 2024).
4.10. Error Handling
@@
Error handling vulnerabilities arise when a system or application fails to properly manage exceptional conditions. Poor error handling can lead to a variety of issues, including security vulnerabilities, system crashes, or unreliable software behavior. One common issue in error handling is the failure to check or handle return values from functions or system calls properly. A few studies focused on error handling vulnerabilities (Sandoval et al., 2023; Elgedawy et al., 2024; Rabbi et al., 2024), however, none of these studies specifically addressed this category of vulnerabilities, which presents another topic for future research. Among the various CWS, CWE-703: Improper Check or Handling of Exceptional Conditions has been specifically highlighted in the literature as a notable issue (Rabbi et al., 2024).
4.11. LLMs in Code Generation: Balancing Risks and Benefits
@@
As we have described throughout this section, LLMs often produce code with security vulnerabilities (Perry et al., 2023; Elgedawy et al., 2024; Siddiq et al., 2024; Tóth et al., 2024; Tihanyi et al., 2024; Fu et al., 2023). Interestingly, when ChatGPT fixes human-written code, the revised version often contains more security vulnerabilities compared to code generated from scratch by ChatGPT (Rabbi et al., 2024).
@@
Despite these challenges, LLMs remain valuable tools in coding tasks. Research indicates that code generated by ChatGPT generally introduces fewer CWE issues compared to code found on Stack Overflow (Hamer et al., 2024). Moreover, LLMs do not significantly increase the incidence rates of severe security bugs (Sandoval et al., 2023), and tools like GitHub Copilot have been shown to produce code with fewer security vulnerabilities than code written by human developers (Asare et al., 2023).
@@
Even though LLMs may generate vulnerable code, the literature suggests that using more effective prompts can guide them to produce safer code (Liu et al., 2024c; Khoury et al., 2023). Through improved re-prompting, the security issues initially present in LLM-generated code can often be addressed, making LLMs valuable tools for creating secure code, when used appropriately. This topic will be explored more in details in Section 7.
5. LLMs for Vulnerability Detection
While LLMs can introduce vulnerabilities into the code they generate, the same models also offer opportunities to improve software security, particularly through their use in vulnerability detection.
@@ 5. LLMs for Vulnerability Detection
Security code reviews performed manually require significant time and effort, especially in large-scale open-source projects with large numbers of contributions. Hence, the automated tools able to identify security vulnerabilities during code review are highly beneficial (Paul et al., 2021). Among these tools, LLMs emerged as a promising solution for detecting security vulnerabilities.
@@
5.1. Uses of LLMs for Vulnerability Detection
@@
The work by Mohajer et al. (Mohajer et al., 2024) demonstrated the ability of ChatGPT to perform static bug detection and false positive warning removal. The evaluation focused on two critical types of bugs often targeted by Static Analysis Tools (SATs): Null Dereference and Resource Leak. For detecting Null Dereference, ChatGPT achieved around 68% accuracy and 64% precision, and for detecting Resource Leak, about 77% accuracy and 83% precision. Concerning false-positive warning removal, ChatGPT reached around 94% precision for Null Dereference bugs and 63% for Resource Leak bugs.
@@
Thapa et al. (Thapa et al., 2022) empirically evaluated the performance of several transformer-based language models for software vulnerability detection. The authors presented a framework that allows for translating C/C++ source code into vectorized inputs, a format suitable for vulnerability analysis. The authors tested various models on software vulnerability datasets, and found that, concerning Buffer Errors and Resource Management Errors, GPT-2 Large and GPT-2 XL achieved the highest F1-score of all the models, with 95.51% and 95.40%, respectively.
@@
A study by Zhang et al. (Zhang et al., 2024d) focused on LLMs capabilities for automated vulnerability detection. They tested over 10 different LLM models using various techniques like zero-shot and one-shot learning (for more details on prompting techniques, we refer the reader to Section 8), and fine-tuning. They applied these models to various datasets of source code, exhibiting C/C++ and smart contract vulnerabilities. The best performing model, CodeLlama-7B, achieved the highest F1-score of 82% with discriminative fine-tuning, along with a precision of approximately 89% and a recall of around 78%.
@@
According to Wen et al. (Wen et al., 2024a), existing pre-trained models for vulnerability detection often miss important details and struggle with complex code. As a consequence, performing accurate vulnerability detection can be difficult. To address this, Wen et al. (Wen et al., 2024a) proposed a Structured Natural Language Comment tree-based Vulnerability Detection framework, based on pre-trained models, named SCALE. This framework is composed of three main modules. The first module, Comment Tree Construction, helps the model understand code semantics, by incorporating LLMs for comment generation and adding the comment node to Abstract Syntax Trees (ASTs). The second module, Structured Natural Language Comment Tree Construction, refines this process by explicitly involving code execution sequences and combining code syntax templates with the comment tree. The third module, SCT-Enhanced Representation, uses these improved comment trees to capture vulnerability patterns and better understand vulnerabilities in the code. Results show that SCALE outperforms existing methods, including supervised, retrained model-based, and other LLM-based approaches in detecting vulnerabilities.
@@
Omar et al. (Omar and Burrell, 2023) proposed a Knowledge Distillation (KD) technique, designed to improve vulnerability detection. In KD, a smaller, simpler model (the “student”) learns to replicate the behavior of a larger, more complex model (the “teacher”). The goal is to achieve similar or even better performance with the smaller model (Beyer et al., 2022). They assessed the performance of the KD when applied to various classifiers, including LLMs such as GPT-2 and CodeBERT, as well as Recurrent Neural Network (RNN) models such as LSTM. They revealed that GPT-2 performed outstandingly, achieving an F1 score of 92.4% on the SARD dataset (Black et al., 2017). While GPT-2 and CodeBERT showcased promising results, it was evident that LSTM-based models still have certain limitations in the context of this application.
@@
Wen et al. (Wen et al., 2024b) highlighted that current methods and datasets focus on detecting intra-procedural vulnerabilities, those found within single functions, overlooking vulnerabilities occurring across multiple functions. To address this, they proposed VulEval, a repository-level evaluation system designed to evaluate both inter- and intra-procedural vulnerabilities. VulEval consists of three phases: the first detects intra-procedural vulnerabilities from code, the second retrieves the most relevant dependencies from call graphs, and the third detects inter-procedural vulnerabilities by incorporating these dependencies. Using VulEval over a large-scale dataset, which includes both intra- and inter-procedural vulnerabilities, they evaluated 19 detection methods: 4 program analysis-based methods, 2 supervised learning-based methods, 7 fine-tuning-based methods, and 6 prompt-based methods. Overall, incorporating vulnerability-related dependencies improved repository-level detection performance, over function-level detection.
5.2. Issues in Using LLMs for Vulnerability Detection